### v1

# Fourier transform of BV functions, isoperimetry, and discrepancy theory

Abstract
This paper investigates the Fourier transform of BV functions. We obtain an averaged Plancherel formula for BV functions and a characterisation of sets of finite perimeter in terms of their Fourier transform. We also sharpen a result of Herz concerning the “set derivative” of the Fourier transform of the characteristic function of a set. Finally, we obtain a new proof of the isoperimetric inequality and sharp bounds on the $L^{2}$-discrepancy of BV functions.
1 Introduction
Fourier analysis has proved to be a very effective tool for studying solutions of partial differential equations and their regularity. Indeed, when the equation has some scaling properties and symmetries, the Fourier transform allows it to pass from a differential problem to an algebraic one, enormously simplifying the calculations. The applications, of course, do not limit themselves to explicit formulae; for example, one can study the regularity of functions knowing the equivalence between Sobolev norms and weighted $L^{2}$-norms of the Fourier transform, namely
$[u]^{2}_{\dot{{\rm H}}^{k}(\mathbb{R}^{d})}=C_{d,k}\int_{\mathbb{R}^{d}}|\xi|^{2k}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi,$
where $[\cdot]_{\dot{{\rm H}}^{k}(\mathbb{R}^{d})}$ with $k\in\mathbb{R}$ is the $H^{k}$-seminorm, and $C_{d,k}$ is a positive constant.
Another remarkable example is the strong correlation between the Fourier restriction conjecture and the decay of solutions of certain wave problems (for example, the Schrödinger problem or the Klein-Gordon equations). Indeed, in the seminal paper [Str77], the author proves, among other things, that estimates on the solutions of such kind of equations are equivalent, if reinterpreted correctly, to estimates on the Fourier transform of functions restricted (or rather, integrated) to a lower dimensional manifold.
In this paper, our approach is the opposite: we employ tools from PDEs and geometric measure theory to establish the precise asymptotic behaviour of integral quantities involving the Fourier transform of BV functions. The applications of the techniques coming from the theory of BV functions and sets of finite perimeter (which is distributional in nature) seem not to be fully explored in the context of Fourier analysis. However, the opposite phenomenon has already occurred. For some examples, one can look at [DPR16] and [ARDPHR19], where the authors are able (broadly speaking) to control the set where a measure is concentrated (and its absolute continuity with respect to the Lebesgue measure $\mathscr{L}^{d}$) by looking in the Fourier space at the “singular” frequencies: as a corollary, for example, they can retrieve the celebrated Alberti Rank One theorem (see [Alb93]).
As mentioned, we shall follow the opposite approach. Further, we plan to refine the techniques of the present work in order to perform a finer analysis on the restriction of the Fourier transform of BV functions and sets of finite perimeter.
Our main result can be stated as follows (see Section 2 for the relevant definitions).
Theorem 1.1.
Let $u,v\in{\rm BV}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, then it holds
$\lim_{R\to+\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}\hat{u}(\xi)\bar{\hat{v}}(\xi)\,\mathrm{d}\xi=\frac{1}{2\pi^{2}}\int_{J_{u}\cap J_{v}}(u^{+}-u^{-})(v^{+}-v^{-})\,\nu_{J_{u}}\cdot\nu_{J_{v}}\,\mathrm{d}\mathcal{H}^{d-1}.$ (1.1)
In particular, if $u=v$, we get
$\lim_{R\to+\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi=\frac{1}{2\pi^{2}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}=:\frac{\mathcal{J}(u)}{2\pi^{2}}.$ (1.2)
Working with characteristic functions of sets, we get the following corollary.
Corollary 1.2.
Let $\Omega_{1},\Omega_{2}\subset\mathbb{R}^{d}$ be sets of finite volume and perimeter, then
$\lim_{R\to\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}\widehat{\mathds{1}}_{\Omega_{1}}(\xi)\bar{\widehat{\mathds{1}}}_{\Omega_{2}}(\xi)\,\mathrm{d}\xi=\frac{1}{2\pi^{2}}\underset{\mathcal{F}\Omega_{1}\cap\mathcal{F}\Omega_{2}}{\int}\nu_{\mathcal{F}\Omega_{1}}\cdot\nu_{\mathcal{F}\Omega_{2}}\,\mathrm{d}\mathcal{H}^{d-1}.$ (1.3)
In particular, if $\Omega_{1}=\Omega_{2}=\Omega$, we have
$\lim_{R\to\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi=\frac{{\rm Per}(\Omega)}{2\pi^{2}}.$ (1.4)
Finally, if
$\limsup_{R\to\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi<+\infty,$ (1.5)
then $\Omega$ is of finite perimeter.
A few comments on both results are in order. For what concerns Theorem 1.1, we retrieve a sort of averaged Plancherel theorem for BV functions: if $u=v$, one can use the standard properties of the Fourier transform to infer that
$\lim_{R\to+\infty}\frac{1}{R}\int_{B_{R}(0)}|\mathcal{F}\{Du\}|^{2}(\xi)\,\mathrm{d}\xi=2\mathcal{J}(u).$ (1.6)
The latter formula is particularly interesting because if $u\in{\rm BV}(\mathbb{R}^{d})$, then the following decomposition for the gradient holds (see also (2.2) below):
$Du=\nabla u\mathscr{L}^{d}+D^{j}u+D^{c}u.$ (1.7)
The first term on the right is the absolutely continuous part, whereas $D^{j}u$ is the jump part, and $D^{c}u$ is the Cantor part. In (1.6), the jump part $D^{j}u$ alone appears in the limit, and this has the following heuristic interpretation: the Fourier transform tells how influent a certain oscillation is, whence when looking at increasingly high frequencies, one expects to observe the faster oscillations. This indicates that jumps in the function, where the oscillation frequency is “infinity”, are the only ones detected in the limit. The second observation we point out is that by integrating by parts, one can deduce a further corollary.
Corollary 1.3.
Let $u\in BV(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, then
$\lim_{R\to\infty}R\int_{B_{R}^{c}(0)}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi=\frac{\mathcal{J}(u)}{2\pi^{2}}.$ (1.8)
The equality in (1.8) is significant for at least two reasons. The first one is that it quantifies the error which is committed by cutting off high frequencies in Plancherel formula; more precisely, we get
$\int_{\mathbb{R}^{d}}|{u}|^{2}(\xi)\,\mathrm{d}\xi=\int_{B_{R}(0)}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi+\frac{\mathcal{J}(u)}{2R}+o(R^{-1}).$
Moreover, considering $u=\mathds{1}_{\Omega}$ in equation (1.8), we have
$\lim_{R\to\infty}R\int_{B_{R}^{c}(0)}|\hat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi=\frac{{\rm Per}(\Omega)}{2\pi^{2}},$ (1.9)
which extends [Mon94, Ch. 6 Thm.3] to any dimension and for any set of finite perimeter (the author proved (1.9) for sets in $\mathbb{R}^{2}$ with piecewise-$\mathcal{C}^{1}$ boundary).
Moving onto Corollary 1.2, we present a newfound characterization of sets of finite perimeter that does not involve distributional calculus and derivatives. Additionally, (1.4) actually improves the estimate obtained in [Her62, Thm. 2], where for every convex set $\Omega$, the author proves the upper bound
$\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi\leq C_{d}{\rm Per}(\Omega),$
for a positive dimensional constant $C_{d}>0$. Further, the author claims that the proof works for general sets without prescribing their regularity, and indeed, its proof can be adjusted for sets of finite perimeter by using standard calculus tools for ${\rm BV}$ functions. We remark that our methods not only improve the upper bound in [Her62, Thm. 2], but rather give a precise asymptotic.
Finally, we mention that in [Str90], the author obtains analogous results for real-valued measures (possibly fractal), asking for some local uniform bounds on the measure of balls. Prior to the latter, in [AH76], the authors study the case of measures supported on $\mathcal{C}^{1}$ submanifolds of $\mathbb{R}^{d}$, obtaining similar estimates.
The second main result we obtain is the characterization of the “set-theoretical derivative” of the Fourier transform of a set of finite perimeter, requiring the so-called outer Minkowski content $\mathcal{S}\mathcal{M}(\Omega)$ (see Definition 2.8 below) to be equal to the total perimeter. Namely, we have the following.
Theorem 1.4.
Let $\Omega\subset\mathbb{R}^{d}$ be a closed set of finite perimeter, then
$\lim_{h\to 0}\frac{\hat{\mathds{1}}_{\Omega+hB_{1}(0)}(\xi)-\hat{\mathds{1}}_{\Omega}(\xi)}{h}=\mathcal{F}\{{|D\mathds{1}_{\Omega}|}\}(\xi)\quad\forall\xi\in\mathbb{R}^{d}$ (1.10)
if and only if ${\rm Per}(\Omega)=\mathcal{S}\mathcal{M}(\Omega)$.
Theorem 1.4 is a refined version of [Her62, Thm. 1]; the author is indeed able to prove the theorem for convex sets, and its proof relies on techniques of differential geometry. The hypotheses under which we can obtain (1.10) are the weakest possible, and, for example, any set with Lipschitz boundary satisfies our assumptions (see [ACV08]). Additionally, our proof happens to be more straightforward since it relies on basic functional analytic techniques and geometric measure theory tools.
As a first application, we give another proof (which, to the authors’ knowledge, appears to be new) of the isoperimetric inequality by combining estimates on the behaviour of the Fourier transform of sets at small and large frequencies (see Theorem 4.2): our inequality is not sharp, however we will investigate this matter in a future work.
Last, we show how the asymptotic relation in equation 1.2 turns out to be key for improving a major result in the theory of irregularities of distribution, also known as discrepancy theory. Namely, we investigate the approximation of the $d$-dimensional Lebesgue measure through samplings by Dirac deltas; this can equivalently be considered as a problem in $\mathbb{R}^{d}$ or in the $d$-dimensional torus $\mathbb{T}^{d}$. In the latter setting, for a set $\Omega\subset\mathbb{T}^{d}$, and for a set of $N$ points $\mathcal{P}_{N}\subset\mathbb{T}^{d}$, the classical discrepancy refers to the quantity
$\mathcal{D}(\Omega;\mathcal{P}_{N})=\sum_{{p}\in\mathcal{P}_{N}}\mathds{1}_{\Omega}({p})-N|\Omega|.$
This field had a pioneering advancement due to [Rot54], where a new geometric point of view was introduced, leading to a wider employment of harmonic analysis techniques. Namely, he showed that there exists a positive absolute constant $C>0$ such that for every $N\geq 1$, it holds
[…]
As later shown in [Dav56], the order $\log N$ is sharp and cannot be improved. On the other hand, considering the wider class of convex bodies rather than just rectangles, one can achieve a polynomial lower bound, as first noticed in [Sch69] when studying the discrepancy of a disk.
A successive major development in the field followed by the work of Beck [Bec87], where the author studied the discrepancy over affine transformation. We introduce an original setting to describe his work. Consider a real function $u\in L^{1}(\mathbb{R}^{d})$, and let $\tau\in\mathbb{R}^{d}$ be a translation factor, let $\delta\geq 0$ be a dilation factor, and let $\rho\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a rotation; in particular, we identify ${\rm SO}(d)$ with the set of all rotations. We define the affine transformation
$[\tau,\delta,\rho]u({x})=u\left(\tau+\rho{x}/\delta\right),$
and from classic properties of the Fourier transform, one has that
$\mathcal{F}\left\{[\mathbf{0},\delta,\rho]u\right\}=\delta^{d}\,[\mathbf{0},\delta^{-1},\rho^{-1}]\mathcal{F}\left\{u\right\}.$ (1.11)
Moreover, we consider the periodization functional ${\mathfrak{P}}\colon L^{1}(\mathbb{R}^{d})\to L^{1}(\mathbb{T}^{d})$, in the sense that
${\mathfrak{P}}\{u\}({x})=\sum_{{n}\in\mathbb{Z}^{d}}u({x}+{n}).$
Then, for a function $u\in L^{1}(\mathbb{R}^{n})$, and for a set of $N$ points $\mathcal{P}_{N}\subset\mathbb{T}^{d}$, we can extend the previous notion of discrepancy by defining
$\mathcal{D}(u;\mathcal{P}_{N})=\sum_{{p}\in\mathcal{P}_{N}}{\mathfrak{P}}\{u\}({p})-N\hat{u}(0).$ (1.12)
In this same setting, we define the $L^{2}$-discrepancy over affine transformations as
[…]
Hence, the main result in [Bec87] can be expressed in the following way: there exists a positive dimensional constant $C_{d}$ such that for every compact and convex set $\Omega\subset\mathbb{R}^{d}$ with non-empty interior, it holds
$\liminf_{N\to+\infty}N^{(1-d)/d}\inf_{\mathcal{P}_{N}\subset\mathbb{T}^{d}}\mathcal{D}_{2}(\mathds{1}_{\Omega};\mathcal{P}_{N})\geq C_{d}{\rm Per}(\Omega).$ (1.14)
The order $N^{(1-d)/d}$ happens to be sharp, as later shown through probabilistic methods in [BC90], although in there, a generic set-depending constant $C_{\Omega}$ replaces the dependence on the perimeter. For a different approach that employs Fourier analysis, the reader can consult [Tra14, Ch. 8], where a well-presented merging of [Ken48] and [Pod91] gives the same upper bound (still, employing a generic constant $C_{\Omega}$ rather than the perimeter).
A few years later, Montgomery [Mon94, Ch. 6] showed that inequality 1.14 holds for every compact set $\Omega\subset\mathbb{R}^{2}$ with piecewise-$\mathcal{C}^{1}$ boundary and non-empty interior, therefore trading convexity for a requirement of regularity. His proof employs Fourier series and a refined version of an argument of Cassels [Cas56] for estimating sums of complex exponentials; the latter is simply described in Section 4, and we point out that a similar argument holds on manifolds (see [BGG21]). More generally, the harmonic analysis techniques introduced turn out to be suitable for investigating various $L^{2}$-discrepancy questions, and we refer to [BT22] and [Ber24] for some recent developments.
It is time to state the sharp bounds we prove in Section 4, starting from the upper one.
Theorem 1.5.
There exists a positive dimensional constant $C_{d}$ such that for every $u\in{\rm BV}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, it holds
$\limsup_{N\to+\infty}N^{(1-d)/d}\inf_{\mathcal{P}_{N}\subset\mathbb{T}^{d}}\mathcal{D}_{2}(u;\mathcal{P}_{N})\leq C_{d}\,\mathcal{J}(u).$
Once equation 1.2 is established, the proof of Theorem 1.5 only employs a sum of uniform lattices. The latter estimate is the best achievable with respect to the order of $N$ and the dependence on $u$, as the following result shows.
Theorem 1.6.
There exists a positive dimensional constant $C_{d}$ such that for every $u\in{\rm BV}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$, it holds
$\liminf_{N\to+\infty}N^{(1-d)/d}\inf_{\mathcal{P}_{N}\subset\mathbb{T}^{d}}\mathcal{D}_{2}(u;\mathcal{P}_{N})\geq C_{d}\,\mathcal{J}(u).$
Again, the proof relies on the asymptotic in equation 1.2 and employs a refinement of the arguments in [Mon94, Ch. 6]. Last, we remark that Theorem 1.6 improves on the estimate of Beck and Montgomery in equation 1.14 since, for instance, no assumption of regularity or convexity is required.
Acknowledgements .
We thank Luca Brandolini, Camillo Brena, Michele Caselli, Leonardo Colzani, Giacomo Gigante, Giorgio Stefani and Giancarlo Travaglini for the useful feedback and discussions.
2 Preliminaries
In this section, we shall introduce the main tools from the theory of PDEs, Fourier analysis and geometric measure theory. For a better account of these topics, the interested reader can consult [Eva98] and [H6̈3] for what concerns PDEs and Fourier analysis. On the other hand, for what concerns geometric measure theory, we advise [AFP00] and [Mag12].
First of all, we define the Fourier transform of an $L^{1}$ function as
$\mathcal{F}\{u\}(\xi)=\hat{u}(\xi)=\int_{\mathbb{R}^{d}}u(x)e^{-2\pi i\xi\cdot x}\,\mathrm{d}x,$
with clear extension to the space of tempered distributions $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, with $\bar{\hat{u}}$ being its complex conjugate. This choice of normalization implies the following version of Plancherel theorem: for $u,v\in L^{2}(\mathbb{R}^{d})$ we have
$\int_{\mathbb{R}^{d}}u(x)v(x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}\hat{u}(\xi)\bar{\hat{v}}(\xi)\,\mathrm{d}\xi.$
Consider now a vector-valued measure $\vec{\mu}=(\mu_{1},...,\mu_{d})$ in $\mathbb{R}^{d}$. We define the total variation of $\vec{\mu}$ as the measure $|\vec{\mu}|$ which to any Borel set $E$ assigns the value
$|\vec{\mu}|(E):=\sup\bigg{\{}\sum_{E_{n}}\|\mu(E_{n})\|_{\mathbb{R}^{d}},\;{\rm with}\;(E_{n})_{n}\;{\rm Borel}\;{\rm partition}\;{\rm of}\;E\bigg{\}}$ (2.1)
We say that a measure $\nu$ (possibly signed and vector-valued) is finite if $|\nu|(\mathbb{R}^{d})<+\infty$ and we shall denote the vector space of finite measures valued in $\mathbb{R}^{k}$ with $\mathcal{M}(\mathbb{R}^{d};\mathbb{R}^{k})$ (if $k=1$ we shall simply write $\mathcal{M}(\mathbb{R}^{d})$): equipping such space with the norm $\|\mu\|_{\rm TV}=|\mu|(\mathbb{R}^{d})$ makes it a Banach space. Then, for a Borel set $E$ and a Borel measure $\mu$, we shall write $\mu_{|E}$ to denote the Borel measure such that $\mu_{|E}(A)=\mu(E\cap A)$ for any Borel set $A$.
We will use the following notion of convergence for sequences of measures.
Definition 2.1 (Weak convergence).
We say that a sequence of measures $(\mu_{n})_{n\in\mathbb{N}}\subset\mathcal{M}(\mathbb{R}^{d})$ weakly converges to a measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ if for every $\varphi\in\mathcal{C}_{b}(\mathbb{R}^{d})$ we have
$\lim_{n\to\infty}\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu_{n}=\int_{\mathbb{R}^{d}}\varphi\,\mathrm{d}\mu.$
We shall now define the class of functions of bounded variation.
Definition 2.2 (BV functions).
We say that $u\in L^{1}(\mathbb{R}^{d})$ is a function of bounded variation, $u\in{\rm BV}(\mathbb{R}^{d})$ for short, if $Du\in\mathcal{M}(\mathbb{R}^{d};\mathbb{R}^{d})$, where $Du$ is the distributional gradient of $u$.
We shall now recall the decomposition of the gradient of a ${\rm BV}$ function: indeed, if $u\in{\rm BV}(\mathbb{R}^{d})$, then $Du$ can be written as (see [AFP00, Thm. 3.78])
$Du=\nabla u\mathscr{L}^{n}+(u^{\vee}-u^{\wedge})\nu_{S_{u}}\mathcal{H}^{d-1}_{|S_{u}}+D^{c}u.$ (2.2)
In the previous decomposition, $\nabla u$ is the density of the absolutely continuous part with respect to $\mathscr{L}^{n}$, while the second term, often denoted with $D^{j}u$, is the jump part of the gradient, where
$u^{\vee}(x)$ $:=\inf\big{\{}t\in[-\infty,\infty]:\;\lim_{\rho\to 0^{+}}\rho^{-d}|\{u>t\}\cap B_{\rho}(x)|=0\big{\}},$
$u^{\wedge}(x)$ $:=\sup\big{\{}t\in[-\infty,\infty]:\;\lim_{\rho\to 0^{+}}\rho^{-d}|\{u<t\}\cap B_{\rho}(x)|=0\big{\}}.$
Moreover $\nu_{S_{u}}$ is the measure theoretic normal and $S_{u}$ is the approximate discontinuity set (see [AFP00, Def. 3.63] ): we have $x\in S_{u}^{c}$ if there exists $a\in\mathbb{R}$ such that
$\lim_{\rho\to 0^{+}}\fint_{B_{\rho}(x)}|u(y)-a|\,\mathrm{d}y=0.$ (2.3)
Finally, $D^{c}u$ is the Cantor part of the measure (again, we refer to [AFP00] for the relevant definitions and properties). We shall use the fact that there exists a Borel set $J_{u}\subseteq S_{u}$ (which is also $\mathcal{H}^{d-1}$-rectifiable) called jump set of $u$, such that $\mathcal{H}^{d-1}(S_{u}\setminus J_{u})=0$ and
$Du=\nabla u\mathscr{L}^{d}+(u^{+}-u^{-})\nu_{J_{u}}\mathcal{H}^{d-1}_{J_{u}}+D^{c}u,$ (2.4)
where
$\lim_{\rho\to 0^{+}}\fint_{B_{\rho}^{+}(x)}|u(y)-u^{+}(x)|\,\mathrm{d}y=\lim_{\rho\to 0^{+}}\fint_{B_{\rho}^{-}(x)}|u(y)-u^{-}(x)|\,\mathrm{d}y=0$ (2.5)
with $B_{\rho}^{+}(x)=\{x\in B_{\rho}(x):\langle x,\nu_{J_{u}}\rangle>0\}$ and $B_{\rho}^{-}(x)=\{x\in B_{\rho}(x):\langle x,\nu_{J_{u}}\rangle<0\}$. For a reference, the interested reader can consult [AFP00] (or [Rin18, p. 280]).
Definition 2.3 (Finite perimeter set).
We say that a Borel set $\Omega\subset\mathbb{R}^{d}$ has finite perimeter, and we write ${\rm Per}(\Omega)<+\infty$, if $\mathds{1}_{\Omega}\in{\rm BV}(\mathbb{R}^{d})$. Finally we define its perimeter as ${\rm Per}(\Omega)=|D\mathds{1}_{\Omega}|(\mathbb{R}^{d})$.
Thanks to the isoperimetric inequality, if ${\rm Per}(\Omega)<+\infty$, then either $\Omega$ or $\Omega^{c}$ has finite measure, whence we shall work assuming that $\Omega$ has finite measure.
We recall that it makes sense to compute the Fourier transform of a measure. Indeed the Fourier transform $\mathcal{F}:\mathcal{M}(\mathbb{R}^{d})\to\mathcal{C}_{b}(\mathbb{R}^{d})$ is a linear and continuous operator with norm equal to $1$ (the same holds componentwise if we consider the operator $\mathcal{F}:\mathcal{M}(\mathbb{R}^{d};\mathbb{R}^{d})\to\mathcal{C}_{b}(\mathbb{R}^{d};\mathbb{R}^{d})$).
For a set of finite perimeter, it is possible to define the notion of reduced boundary, which amounts to identifying the more “regular” points of the boundary.
Definition 2.4.
We define the Borel set $\mathcal{F}\Omega$ and we call it reduced boundary of $\Omega$ as the set of points $x\in{\rm supp}|D\mathds{1}_{\Omega}|$ such that the limit
$\nu_{\Omega}(x):=\lim_{r\to 0^{+}}\frac{D\mathds{1}_{\Omega}(B_{r}(x))}{|D\mathds{1}_{\Omega}|(B_{r}(x))}$ (2.6)
exists and equals $1$.
Remark 2.5.
Observe that $\mathcal{F}\Omega$ is always a subset of the topological boundary, and it may happen that the two do not coincide. Indeed, the reduced boundary of a square consists of all topological boundaries except the $4$ points in the angles.
We are now in the position to recall part of the celebrated structural result for sets of finite perimeter, namely De Giorgi theorem (see [AFP00, Thm. 3.59])
Theorem 2.6 (De Giorgi).
Let $\Omega\subset\mathbb{R}^{d}$ be a set of finite perimeter, then $\mathcal{F}\Omega$ is countably $\mathcal{H}^{d-1}$ rectifiable and $|D\mathds{1}_{\Omega}|=\mathcal{H}^{d-1}_{|\mathcal{F}\Omega}$.
Remark 2.7.
We shall not delve into the definition of rectifiability in the above sense (again, we refer the reader to [AFP00] or [Mag12]): one can think of it as a measure-theoretic notion of being $(d-1)$-dimensional.
We shall write $\Omega_{r}$ for the following set
$\Omega_{r}:=\{x\in\mathbb{R}^{d}:\;{\rm d}(\Omega,x)\leq r\},$
which coincides with the Minkowski sum $\Omega+B_{r}(0)$. We have the following definition
Definition 2.8 (outer Minkowski content).
We say that a Borel set $\Omega\subseteq\mathbb{R}^{d}$ has a finite outer Minkowski content if the limit
$\mathcal{S}\mathcal{M}(\Omega):=\lim_{r\to 0^{+}}\frac{|\Omega_{r}\setminus\Omega|}{r}$
exists and is finite.
We now recall an integral Abelian-Tauberian theorem. It is usually stated as two different theorems, but we shall write it as a single theorem; for the proof, we refer to [Fel71, p. 443 Thm. 1] (see also [AHT18, Appendix]).
Theorem 2.9 (Abelian-Tauberian).
Let $\nu$ be a $\sigma$-finite positive Borel measure and let $\gamma\in[0,\infty)$, then
$\lim_{t\to 0^{+}}t^{\gamma}\int_{0}^{\infty}e^{-t\lambda}\,\mathrm{d}\nu(\lambda)=C\qquad\iff\qquad\lim_{a\to\infty}\frac{\nu([0,a])}{a^{\gamma}}=\frac{C}{\Gamma(\gamma+1)}.$
Now, following the notation in [MPPP07], for $u\in L^{1}$, we denote with $T(t)u$ the unique solution of
$\begin{cases}\partial_{t}v=\Delta v\qquad{\rm in}\;\mathbb{R}^{d}\times(0,\infty)\\ v(0,x)=u\qquad{\rm on}\;\mathbb{R}^{d},\end{cases}$ (2.7)
evaluated at time $t>0$. Recall, moreover, that one has the explicit representation via convolution
$(T(t)u)(x)=\int_{\mathbb{R}^{n}}H_{\mathbb{R}^{d}}(x-y,t)u(y)\,\mathrm{d}y,$
where $H_{\mathbb{R}^{d}}$ is the classical Gaussian heat kernel
$H_{\mathbb{R}^{d}}(z,t)=\frac{1}{(4\pi t)^{\frac{d}{2}}}e^{-\frac{|z|^{2}}{4t}}.$ (2.8)
We now recall the result in [MPPP07, Thm. 4.3] (see also [GS24] for a recent generalization to various differential operators).
Theorem 2.10.
Let $u,v\in{\rm BV}(\mathbb{R}^{d})\cap L^{2}(\mathbb{R}^{d})$, then
$\lim_{t\to 0^{+}}\frac{\sqrt{\pi}}{\sqrt{t}}\langle u-T(t)u,v\rangle=\int_{J_{u}\cap J_{v}}(u^{+}-u^{-})(v^{+}-v^{-})\nu_{J_{u}}\cdot\nu_{J_{v}}\,\mathrm{d}\mathcal{H}^{d-1}.$ (2.9)
Remark 2.11.
For how the statement is written, we are not able to take advantage of it: indeed, in this way, we can only prove that the map
$H(t)=\int_{\mathbb{R}^{d}}(T(t^{2})u)(x)v(x)\,\mathrm{d}x$ (2.10)
is differentiable at $t=0$. What we need for our applications is that $\lim_{t\to 0^{+}}H^{\prime}(t)$ exists (and in this case, it clearly coincides with (2.9)). In [MPPP07], the authors are able to prove that such a limit exists for characteristic functions of sets. However, they don’t do it for general BV functions. To obtain (2.9), they rely on the result for characteristic functions and use Coarea formula, therefore bypassing the existence of $\lim_{t\to 0^{+}}H^{\prime}(t)$: this is the content of Proposition 3.2.
Finally we recall [MPPP07, Thm 3.4]
Theorem 2.12.
Let $\Omega\subset\mathbb{R}^{d}$ be such that either $|\Omega|<+\infty$ or $|\Omega^{c}|<+\infty$. If
$\liminf_{t\to 0^{+}}\frac{1}{t}\int_{\Omega^{c}}T(t^{2})\mathds{1}_{\Omega}(x)\,\mathrm{d}x<+\infty,$
then $\Omega$ has finite perimeter.
3 Proof of the main results
The following result is essential to obtain Theorem 1.4, and it is a refinement of [ACV08, Prop. 2].
Proposition 3.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a closed set. Then $\mathcal{S}\mathcal{M}(\Omega)={\rm Per}(\Omega)$ if and only if the measures
$\mu_{h}:=\frac{\mathds{1}_{\Omega+hB_{1}(0)}-\mathds{1}_{\Omega}}{h}$ (3.1)
weakly converge to $|D\mathds{1}_{\Omega}|$.
Proof.
Let us first assume that $\mathcal{S}\mathcal{M}(\Omega)={\rm Per}(\Omega)<+\infty$. To prove the statement it is enough to check that $\liminf_{h\to 0^{+}}\mu_{h}(U)\geq|D\mathds{1}_{\Omega}|(U)$ for every $U$ open set thanks to Portmanteau theorem (see [Bog07, Thm. 8.2.3]), combined with the fact that $\lim_{h\to 0^{+}}\mu_{h}(\mathbb{R}^{d})={\rm Per}(\Omega)$ by assumption. Therefore, pick the open set $U$ and consider the distance function from $\Omega$, namely the function ${\rm d}_{\Omega}(x):=\inf\{|x-y|:\;y\in\Omega\}$. It is clear that such a function is Lipschitz continuous, and it can be proved that $|\nabla{\rm d}_{\Omega}|=1$ $\mathscr{L}^{d}$-almost everywhere in $\Omega^{c}$ (see [AD00, p. 11]). Given the previous observations, we can apply Coarea formula to get
$\mu_{h}(U)=\frac{1}{h}\underset{\{0<\,\mathrm{d}_{\Omega}<h\}}{\int}\mathds{1}_{U}|\nabla\,\mathrm{d}_{\Omega}|\,\mathrm{d}\mathcal{H}^{d}=\frac{1}{h}\int_{0}^{h}{\rm Per}(\{{\rm d}_{\Omega}>t\};U)\,\mathrm{d}t=\int_{0}^{1}{\rm Per}(\{{\rm d}_{\Omega}>hr\};U)\,\mathrm{d}r.$
Since the sets $\{{\rm d}_{\Omega}>th\}$ converge locally in measure to $\{{\rm d}_{\Omega}>0\}=\Omega^{c}$ as $h\to 0^{+}$, by the lower semicontinuity of the perimeter (see [AFP00, Prop. 3.38 (b)]) and Fatou lemma, we get
$\liminf_{h\to 0^{+}}\mu_{h}(U)\geq{\rm Per}(\Omega^{c};U)={\rm Per}(\Omega;U)=|D\mathds{1}_{\Omega}|(U),$
therefore proving the result.
Conversely let $\mu_{h}$ weakly converge to $|D\mathds{1}_{\Omega}|$: this clearly implies the boundedness of the sequence and moreover $\mu_{h}(\mathbb{R}^{d})\to|D\mathds{1}_{\Omega}|(\mathbb{R}^{d})={\rm Per}(\Omega)$, which is the thesis. ∎
The second result we shall use is an improvement of Theorem 2.10, showing not only that $t\mapsto H$ is differentiable at zero, but also that $\lim_{t\to 0^{+}}H^{\prime}(t)$ exists.
Proposition 3.2.
Let $u,v\in{\rm BV}(\mathbb{R}^{d})$, with $u\in L^{\infty}(\mathbb{R}^{d})$ or $v\in L^{\infty}(\mathbb{R}^{d})$, then
$\lim_{t\to 0^{+}}H^{\prime}(t)=-\frac{1}{\sqrt{\pi}}\int_{J_{u}\cap J_{v}}(u^{+}-u^{-})(v^{+}-v^{-})\nu_{J_{u}}\cdot\nu_{J_{v}}\,\mathrm{d}\mathcal{H}^{d-1}.$ (3.2)
Proof.
Without loss of generality, let us assume $u\in L^{\infty}(\mathbb{R}^{d})$. Clearly $H\in\mathcal{C}^{1}((0,\infty))$ and the following calculations are justified
$H^{\prime}(t)=2t\int_{\mathbb{R}^{d}}\Delta(T(t^{2})u)(x)v(x)\,\mathrm{d}x=-2t\int_{\mathbb{R}^{d}}\nabla(T(t^{2})u)(x)\,\mathrm{d}Dv(x).$ (3.3)
We shall now compute the term $\nabla(T(t^{2})u)(x)$ explicitly. Namely, we have
$\nabla(T(t^{2})u)(x)$ $=\int_{\mathbb{R}^{d}}\nabla_{x}H_{\mathbb{R}^{d}}(x-y,t^{2})u(y)\,\mathrm{d}y$
$=-\frac{1}{2t^{2}}\int_{\mathbb{R}^{d}}(x-y)\frac{e^{-\frac{|x-y|^{2}}{4t^{2}}}}{(4\pi t^{2})^{\frac{d}{2}}}u(y)\,\mathrm{d}y$
$=-\frac{1}{(4\pi)^{\frac{d}{2}}2t}\int_{\mathbb{R}^{d}}ze^{-\frac{|z|^{2}}{4}}u(x-tz)\,\mathrm{d}z.$
Therefore, (3.3) becomes
$H^{\prime}(t)=-\frac{1}{(4\pi)^{\frac{d}{2}}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}ze^{-\frac{|z|^{2}}{4}}u(x-tz)\,\mathrm{d}z\,\mathrm{d}Dv(x).$
Let us now call
$g(x,t):=\int_{\mathbb{R}^{d}}ze^{-\frac{|z|^{2}}{4}}u(x-tz)\,\mathrm{d}z$ (3.4)
and split the integral as follows
$H^{\prime}(t)=-\frac{1}{(4\pi)^{\frac{d}{2}}}\int_{J_{u}}g(x,t)\,\mathrm{d}Dv(x)-\frac{1}{(4\pi)^{\frac{d}{2}}}\int_{S_{u}^{c}}g(x,t)\,\mathrm{d}Dv(x)=:G_{1}(t)+G_{2}(t),$
where $S_{u}$ is the approximate discontinuity set of $u$ (see (2.2)), and $J_{u}$ is the jump set (we are allowed to write the latter decomposition since $\mathcal{H}^{d-1}(S_{u}\setminus J_{u})=0$). Observe, moreover, that $g(\cdot,t)$ is a continuous function for which we have
$\sup_{x\in\mathbb{R}^{d}}|g(x,t)|\leq\|u\|_{\infty}\int_{\mathbb{R}^{d}}|z|e^{-\frac{|z|^{2}}{4}}\,\mathrm{d}z\leq C_{d}\|u\|_{\infty}.$ (3.5)
Claim: $\lim_{t\to 0^{+}}G_{2}(t)=0$.
To prove the claim, observe that one can write
$g(x,t)=\int_{\mathbb{R}^{d}}ze^{-\frac{|z|^{2}}{4}}\big{(}u(x-tz)-\tilde{u}(x)\big{)}\,\mathrm{d}z,$
where $\tilde{u}(x)$ is the approximate limit of $u$ at $x$. Now fix $R>0$ and estimate $g$ as follows
$|g(x,t)|$ $\leq\int_{B_{R}(0)}|z|e^{-\frac{|z|^{2}}{4}}|u(x-tz)-\tilde{u}(x)|\,\mathrm{d}z+\int_{B_{R}^{c}(0)}|z|e^{-\frac{|z|^{2}}{4}}|u(x-tz)-\tilde{u}(x)|\,\mathrm{d}z$
$\leq\int_{B_{R}(0)}|z|e^{-\frac{|z|^{2}}{4}}|u(x-tz)-\tilde{u}(x)|\,\mathrm{d}z+2\|u\|_{L^{\infty}}\int_{B_{R}^{c}(0)}|z|e^{-\frac{|z|^{2}}{4}}\,\mathrm{d}z.$
[…]
$\limsup_{t\to 0^{+}}|g(x,t)|\leq 2\|u\|_{L^{\infty}}\int_{B_{R}^{c}(0)}ze^{-\frac{|z|^{2}}{4}}\,\mathrm{d}z,$
[…]
We are thus left to study what happens in the set $J_{u}$: first of all, we have that $J_{u}$ is countably $\mathcal{H}^{d-1}$-rectifiable, whence we can apply [AFP00, Prop. 3.92 (c)] to deduce that $|D^{c}v|(J_{u})=0$. Moreover, the set $J_{u}$ has Lebesgue measure zero, so we are left with
$G_{1}(t)=-\frac{1}{(4\pi)^{\frac{d}{2}}}\int_{J_{u}\cap J_{v}}g(x,t)\cdot\nu_{J_{v}}(v^{+}-v^{-})\,\mathrm{d}\mathcal{H}^{d-1}(x).$ (3.6)
Now define $f(x,t):=\langle g(x,t),\nu_{J_{v}}(x)\rangle$ and, exploiting the symmetries of the integrand, we write
$f(x,t)$ $=\int_{H_{\nu_{J_{u}}}^{+}}\langle z,\nu_{J_{v}}(x)\rangle e^{-\frac{|z|^{2}}{4}}(u(x-tz)-u^{+}(x))\,\mathrm{d}z-u^{+}(x)\int_{H_{\nu_{J_{u}}}^{-}}\langle z,\nu_{J_{v}}(x)\rangle e^{-\frac{|z|^{2}}{4}}\,\mathrm{d}z$
$+\int_{H_{\nu_{J_{u}}}^{-}}\langle z,\nu_{J_{v}}(x)\rangle e^{-\frac{|z|^{2}}{4}}(u(x-tz)-u^{-}(x))\,\mathrm{d}z-u^{-}(x)\int_{H_{\nu_{J_{u}}}^{-}}\langle z,\nu_{J_{v}}(x)\rangle e^{-\frac{|z|^{2}}{4}}\,\mathrm{d}z$
$=:A+B+C+D,$
where $H^{\underset{-}{+}}_{\nu_{J_{u}}(x)}=\{z\in\mathbb{R}^{d}:\;\langle z,\nu_{J_{u}}(x)\rangle\underset{<}{\geq}0\}$.
Claim: $\lim_{t\to 0^{+}}A=\lim_{t\to 0^{+}}C=0$ for every $x\in J_{u}\cap J_{v}$.
Let us consider the term $A$ (the term $C$ can be treated analogously): we have
$A=\int_{B_{R}\cap H_{\nu_{J_{u}}}^{+}}(\dots)+\int_{B_{R}^{c}\cap H_{\nu_{J_{u}}}^{+}}(\dots).$
The first integral on the previous r.h.s. goes to zero by definition of $u^{+}(x)$, while the second one is equibounded by $\|u\|_{\infty}\int_{B_{R}^{c}(0)}|z|e^{-|z|^{2}/4}\,\mathrm{d}z$, and therefore goes to 00 as $R\to\infty$, proving the claim.
Now observe that since for $\mathcal{H}^{d-1}$-a.e. $x\in J_{u}\cap J_{v}$ we have that $\nu_{J_{v}}(x)=\nu_{J_{u}}(x)\langle\nu_{J_{u}}(x),\nu_{J_{v}}(x)\rangle$, then we get
$B=\frac{(4\pi)^{\frac{d}{2}}}{\sqrt{\pi}}u^{+}(x)\langle\nu_{J_{u}}(x),\nu_{J_{v}}(x)\rangle,\quad D=-\frac{(4\pi)^{\frac{d}{2}}}{\sqrt{\pi}}u^{-}(x)\langle\nu_{J_{u}}(x),\nu_{J_{v}}(x)\rangle.$
A straightforward application of dominated convergence finally gives
$\lim_{t\to 0^{+}}G_{1}(t)=-\frac{1}{\sqrt{\pi}}\int_{J_{u}\cap J_{v}}(u^{+}-u^{-})(v^{+}-v^{-})\nu_{J_{u}}\cdot\nu_{J_{v}}\,\mathrm{d}\mathcal{H}^{d-1},$
which is the thesis. ∎
We now have all the tools to prove Theorem 1.1.
Proof of Theorem 1.1.
We shall actually prove first (1.2). Let $u$ as in the assumptions and consider the quantity
$H(t):=\int_{\mathbb{R}^{d}}(T(t^{2})u)(x)u(x)\,\mathrm{d}x.$
Thanks to Plancherel theorem and the explicit expression of the heat kernel (see (2.8)), we get
$H(t)=\int_{\mathbb{R}^{d}}\mathcal{F}\{T(t^{2})u\}(\xi)\overline{\mathcal{F}\{u\}}(\xi)\,\mathrm{d}\xi=\int_{\mathbb{R}^{d}}e^{-4\pi^{2}|\xi|^{2}t^{2}}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi.$ (3.7)
Observe that $t\mapsto H(t)$ is continuous in $[0,1]$, moreover by Plancherel theorem and dominated convergence $H(0)=\|u\|_{L^{2}}$. Again, because of dominated convergence, it follows that
$H^{\prime}(t)=-8\pi^{2}t\int_{\mathbb{R}^{d}}|\xi|^{2}e^{-4\pi^{2}|\xi|^{2}t^{2}}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi,$ (3.8)
and that $H^{\prime}$ is continuous. Thanks to Proposition 2.9, we get
$t\int_{\mathbb{R}^{d}}|\xi|^{2}e^{-4\pi^{2}|\xi|^{2}t^{2}}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi\to\frac{1}{8\pi^{\frac{5}{2}}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}$
as $t\to 0^{+}$. Now let us pass to polar coordinates to get
$t\int_{0}^{\infty}e^{-4\pi^{2}r^{2}t^{2}}r^{n+1}g(r)\,\mathrm{d}r\sim\frac{1}{8\pi^{\frac{5}{2}}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1},$
where
$g(r):=\int_{\mathbb{S}^{d-1}}|\hat{u}|^{2}(rv)\,\mathrm{d}\sigma(v).$
Let us call $\delta=4\pi^{2}t^{2}$, and then change variables as $r^{2}=s$ to get
$\int_{0}^{\infty}e^{-\delta s}s^{\frac{n}{2}}g(\sqrt{s})\,\mathrm{d}s\sim\frac{1}{2\pi^{\frac{3}{2}}\sqrt{\delta}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}.$ (3.9)
We now set
$\Phi(s):=\int_{0}^{s}z^{\frac{n}{2}}g(\sqrt{z})\,\mathrm{d}z=2\int_{0}^{\sqrt{s}}k^{n+1}g(k)\,\mathrm{d}k=2\int_{B_{\sqrt{s}}(0)}|\xi|^{2}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi.$
With the previous definition, we can write (3.9) as
$\int_{0}^{\infty}e^{-\delta s}\,\mathrm{d}\Phi(s)\sim\frac{1}{2\pi^{\frac{3}{2}}\sqrt{\delta}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}$
when $\delta\to 0$.
Finally, we can apply Theorem 2.9 to deduce that
$\Phi(R)\sim\frac{\sqrt{R}}{\pi^{2}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1},$
as $R\to\infty$, which means that (using $R^{2}$ in place of $R$)
$\lim_{R\to\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}|\hat{u}|^{2}(\xi)\,\mathrm{d}\xi=\frac{1}{2\pi^{2}}\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}.$ (3.10)
To prove (1.1), we shall proceed by polarization: let again $u,v$ as in the assumptions and let us exploit the fact that
$\Re\{\hat{u}(\xi)\bar{\hat{v}}(\xi)\}=\frac{|\widehat{(u+v)}(\xi)|^{2}}{2}-\frac{|\hat{u}(\xi)|^{2}}{2}-\frac{|\hat{v}(\xi)|^{2}}{2},$
the symmetries of the Fourier transform and (1.2) to obtain
$\lim_{R\to\infty}\frac{1}{R}\int_{B_{R}(0)}|\xi|^{2}\hat{u}(\xi)\bar{\hat{v}}(\xi)\,\mathrm{d}\xi$ $=\frac{1}{4\pi^{2}}\bigg{(}\int_{J_{u+v}}((u+v)^{+}-(u+v)^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}$
$-\int_{J_{u}}(u^{+}-u^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}-\int_{J_{v}}(v^{+}-v^{-})^{2}\,\mathrm{d}\mathcal{H}^{d-1}\bigg{)}.$
Now we write
$J_{u+v}$ $=(J_{u}\setminus J_{v})\cup(J_{v}\setminus J_{u})\cup(J_{u}\cap J_{v}).$
$J_{u}$ $=(J_{u}\setminus J_{v})\cup(J_{u}\cap J_{v}),$
$J_{v}$ $=(J_{v}\setminus J_{u})\cup(J_{u}\cap J_{v})$
and we split the integrals on the r.h.s. observing that if $x\in J_{u}\cap J_{v}$ then $(u+v)^{\underset{-}{+}}(x)=u^{\underset{-}{+}}(x)+v^{\underset{-}{+}}(x)$. Finally, easy calculations give (1.1). ∎
Proof of Corollary 1.2.
Thanks to (3.10), we easily recover Montgomery’s result in any dimension and for any set of finite perimeter (and volume). Indeed pick $u=v=\mathds{1}_{\Omega}$. Then Theorem 1.1 gives (1.4).
[…]
$\limsup_{\delta\to 0^{+}}\sqrt{\delta}\int_{0}^{\infty}e^{-\delta s}\,\mathrm{d}\Phi(s)<+\infty,$
with
$\Phi(s)=\int_{0}^{s}z^{\frac{n}{2}}g(\sqrt{z})\,\mathrm{d}z,\quad g(r)=\int_{\mathbb{S}^{d-1}}|\hat{\mathds{1}}_{\Omega}|^{2}(rv)\,\mathrm{d}v.$
Changing variables and using Coarea formula then gives
$\limsup_{t\to 0^{+}}t\int_{\mathbb{R}^{d}}e^{-|\xi|^{2}t^{2}}|\xi|^{2}|\hat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi<+\infty.$
The latter implies, via Plancherel theorem, that
$\limsup_{t\to 0^{+}}|H^{\prime}(t)|<+\infty,$
where $H^{\prime}$ is (3.3) with $u=v=\mathds{1}_{\Omega}$. However, by the mean value theorem and monotonicity properties of $H$, we get
$0\leq\limsup_{t\to 0}\frac{H(0)-H(t)}{t}<+\infty.$ (3.11)
Exploiting the definition of $H$, we can turn (3.11) into
$\limsup_{t\to 0^{+}}\frac{1}{t}\int_{\Omega^{c}}\big{(}T(t^{2})\mathds{1}_{\Omega}\big{)}(x)\,\mathrm{d}x<+\infty,$
and thanks to Theorem 2.12, we can conclude. ∎
Proof of Corollary 1.3.
Before performing the limit in (1.2) by $R^{-3}$ on both sides and integrate from $L>0$ and $\infty$ to get (let us call $C=\mathcal{J}(u)/(2\pi^{2})$)
$\int_{L}^{\infty}R^{-3}\int_{B_{R}(0)}|\xi|^{2}|\widehat{u}|^{2}(\xi)\,\mathrm{d}\xi\sim\frac{C}{L}$
Now integrate by parts the previous expression on the left-hand side to get (use (3.10) to justify what follows)
$\frac{1}{2L^{2}}\int_{B_{L}(0)}|\xi|^{2}|\widehat{u}|^{2}(\xi)|^{2}(\xi)\,\mathrm{d}\xi+\frac{1}{2}\int_{B_{L}^{c}(0)}|\widehat{u}|^{2}(\xi)|^{2}(\xi)\,\mathrm{d}\xi\sim\frac{C}{L}.$
Finally, applying again (1.4), we end up with
$\int_{B_{L}^{c}(0)}|\widehat{u}|^{2}(\xi)|^{2}(\xi)\,\mathrm{d}\xi\sim\frac{C}{L}.$
which is (1.8). Taking $u=\mathds{1}_{\Omega}$ clearly gives (1.9). ∎
We now prove the last result of this section.
Proof of Theorem 1.4.
The proof is an easy corollary of Proposition 3.1; indeed, since $e^{-2\pi i\xi\cdot x}$ is a bounded and continuous function for all $\xi\in\mathbb{R}^{d}$ and the measures $\mu_{h}$ converge in duality with $\mathcal{C}_{b}(\mathbb{R}^{d})$ to $|D\mathds{1}_{\Omega}|$, we have
$\lim_{h\to 0^{+}}\int_{\mathbb{R}^{d}}e^{-2\pi i\xi\cdot x}\,\mathrm{d}\mu_{h}(x)=\int_{\mathbb{R}^{d}}e^{-2\pi i\xi\cdot x}\,\mathrm{d}|D\mathds{1}_{\Omega}|(x),$
which is equivalent to (1.10). The converse is trivial in light of Proposition 3.1 and the fact that the weak convergence always implies the convergence of the masses (or, in other words, that $\lim_{h\to 0^{+}}\mu_{h}(\mathbb{R}^{d})=|D\mathds{1}_{\Omega}|(\mathbb{R}^{d})$). ∎
4 Applications to isoperimetry and discrepancy on BV functions
In this section, we shall show some applications of our results and techniques, the first one being the proof of the celebrated isoperimetric inequality, with a non-sharp constant. First, we shall state and prove an easy consequence of [Her62, Thm. 2].
Lemma 4.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a set of finite volume and perimeter, then there exists a dimensional constant $C_{d}>0$ such that
$R\int_{B_{R}^{c}(0)}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi\leq C_{d}{\rm Per}(\Omega).$ (4.1)
Proof.
From [Her62, Thm. 2] there exists $C_{d}>0$ such that
$\int_{B_{T}(0)}|\xi|^{2}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi\leq C_{d}{\rm Per}(\Omega)T$
for all $T>0$. As in Corollary 1.2, multiply both sides of the previous inequality by $T^{-3}$ and integrate from $R$ to $\infty$. After an integration by parts, we get (4.1). ∎
We are now ready to state and prove the isoperimetric inequality.
Theorem 4.2 (Isoperimetric inequality).
Let $\Omega\subset\mathbb{R}^{d}$ be such that $|\Omega|<+\infty$, then there exists $C_{d}>0$ such that
$|\Omega|\leq C_{d}{\rm Per}^{\frac{d}{d-1}}(\Omega).$ (4.2)
Proof.
Without relabeling, we will write $C_{d}>0$ for any dimensional constant appearing in the proof. Without loss of generality, we can assume that ${\rm Per}(\Omega)<+\infty$. By using Plancherel theorem, we write
$|\Omega|=\int_{B_{R}(0)}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi+\int_{B_{R}^{c}(0)}|\widehat{\mathds{1}}_{\Omega}|^{2}(\xi)\,\mathrm{d}\xi.$
By the fact that $\|\widehat{\mathds{1}}_{\Omega}\|_{\infty}\leq|\Omega|$ and (4.1), we get
$|\Omega|\leq|\Omega|^{2}\omega_{d}R^{d}+C_{d}\frac{{\rm Per}(\Omega)}{R}$
for all $R>0$. Now, we can optimize in $R$, so that by choosing
$R=\bigg{(}\frac{C_{d}{\rm Per}(\Omega)}{d\omega_{d}|\Omega|^{2}}\bigg{)}^{\frac{1}{d+1}},$
we get
$|\Omega|\leq C_{d}|\Omega|^{\frac{2}{d+1}}{\rm Per}^{\frac{d}{d+1}}(\Omega).$
After minor rearrangements, we end up with
$|\Omega|\leq C_{d}{\rm Per}^{\frac{d}{d-1}}(\Omega),$
which is the thesis. ∎
Remark 4.3.
Observe that the constant we obtain, although explicit, is not the sharp one. We plan to refine the inequalities involved in the present proof and refine our analysis in order to obtain the sharp constant, and investigate possible quantitative versions (in the spirit of [FMP08]).
We now turn our attention to $L^{2}$-discrepancy and apply our techniques to prove Theorem 1.5 and Theorem 1.6.
Let us start by exploiting the convolutional structure of (1.12). Namely, if we let $\mu_{\rm L}$ be the Lebesgue measure on $\mathbb{T}^{d}$, and for ${p}\in\mathbb{T}^{d}$ we consider $\mu_{\rm D}({p})$ to be the Dirac delta centered at ${p}$, then by setting
$\tilde{\mu}=\sum_{{p}\in\mathcal{P}_{N}}\mu_{\rm D}(-{p})-N\mu_{\rm L},$
for a function $u\in L^{1}(\mathbb{R}^{d})$, we may rewrite (1.12) as
[…]
Now, for $f\in L^{1}(\mathbb{T}^{d})$ or $f\in\mathcal{M}(\mathbb{T}^{d})$, let ~$\tilde{\mathcal{F}}\{f\}\colon\mathbb{Z}^{d}\to\mathbb{C}$ be the function of the Fourier coefficients of $f$. Then, it is not difficult to see that for $u\in L^{1}(\mathbb{R}^{d})$, and for every ${n}\in\mathbb{Z}^{d}$, it holds
~$\tilde{\mathcal{F}}\circ\mathfrak{P}\{u\}({n})=\mathcal{F}\{u\}({n}).$
Therefore, by applying Plancherel theorem on $\mathbb{T}^{d}$, we get
[…]
where, for the sake of notation, we have set $\mathbb{Z}^{d}_{*}=\mathbb{Z}^{d}\setminus\{\mathbf{0}\}$. Finally, by (1.11) and the symmetry of ${\rm SO}(d)$, we may rewrite (1.13) as
$\mathcal{D}_{2}(u;\mathcal{P}_{N})=\sum_{{n}\in\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{{p}\in\mathcal{P}_{N}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}\int_{{\rm SO}(d)}\int_{0}^{1}\delta^{2d}\left|[\mathbf{0},\delta^{-1},\rho]\mathcal{F}\{u\}({n})\right|^{2}\,\,\mathrm{d}\delta\,\,\mathrm{d}\rho,$
and further, by a change of variable, we get
$\begin{split}\mathcal{D}_{2}(u;\mathcal{P}_{N})=\sum_{{n}\in\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{{p}\in\mathcal{P}_{N}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi.\end{split}$ (4.3)
The rest of the section relies on the following asymptotic relation resulting from integrating by parts in (1.2). Namely, it holds
$\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\sim\frac{{|{n}|^{d}\,\mathcal{J}}(u)}{2d\pi^{2}}.$ (4.4)
Proof of Theorem 1.5.
We will write $c_{i}=c_{i}(d)$ with $i\in\mathbb{N}$ to intend generic dimensional constants throughout the proof. By (4.4), there exists $n_{0}\in\mathbb{N}$ be such that for every ${n}\geq n_{0}$, it holds
$\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\leq\frac{{|{n}|^{d}\,\mathcal{J}}(u)}{d\pi^{2}}.$ (4.5)
First, we prove the upper bound for all positive integers $N$ such that $N^{1/d}\geq n_{0}$ and $N^{1/d}\in\mathbb{N}$. Consider the set of $N$ points $\mathcal{P}_{N}\subset\mathbb{T}^{d}$ to be the periodization of
$\frac{[0,N^{1/d})^{d}\cap\mathbb{Z}^{d}}{N^{1/d}}\subset[0,1)^{d}.$ (4.6)
Hence, it is not difficult to see that
$\sum_{{p}\in\mathcal{P}_{N}}e^{2\pi i{p}\cdot{n}}=\begin{cases}N&{\rm if}\quad{n}\in N^{1/d}\,\mathbb{Z}^{d}\\ 0&{\rm else}\end{cases},$
and therefore, by (4.3) and the assumptions on $N$, we get
$\begin{split}\mathcal{D}_{2}(u;\mathcal{P}_{N})&=\sum_{{m}\in\mathbb{Z}_{*}^{d}}N^{2}|N^{1/d}\,{m}|^{-2d-1}\int_{|\xi|\leq N^{1/d}\,|{m}|}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\\ &\leq\frac{1}{d\pi^{2}}\,\mathcal{J}(u)\,N^{2}\sum_{{n}\in\mathbb{Z}_{*}^{d}}|N^{1/d}\,{m}|^{-d-1}\\ &\leq c_{1}\,\mathcal{J}(u)\,N^{(d-1)/d}.\end{split}$
We now prove the upper bound for every positive integer $N$. As $k$ goes through the positive integers, we consider the recursive decomposition of $N$ defined as
$N_{k}=\max\bigg{\{}n\in\mathbb{N}\,\colon\,n^{1/d}\in\mathbb{N},\>n\leq N-\sum_{j=1}^{k-1}N_{j}\bigg{\}},$
where improper sums are zeros. By the latter definition, we get that
$N-\sum_{j=1}^{k-1}N_{j}\leq(N_{k}^{1/d}+1)^{d}\leq N_{k}+c_{2}N_{k}^{(d-1)/d},$
and therefore, for every integer $k\geq 1$, it holds
$N_{k+1}\leq N-\sum_{j=1}^{k}N_{j}\leq c_{2}N_{k}^{(d-1)/d}.$ (4.7)
Hence, since by definition $N_{1}\leq N$, then we recursively get
$N_{k+1}^{1/d}\leq c_{2}^{1/d}N_{k}^{(d-1)/d^{2}}\leq c_{2}^{(k-1)/d}N^{(d-1)^{k-1}/{d^{k}}}.$ (4.8)
Now, set $K=K(d)$ to be the smallest integer such that
$\frac{(d-1)^{K}}{d^{K}}\leq\frac{d-1}{4d},$
and in particular, notice that by (4.8), we have
$N_{K+1}^{(d-1)/d}\leq c_{3}\,N^{(d-1)/(4d)}.$ (4.9)
Now, consider the remainder
$N_{0}=N-\sum_{j=1}^{K+1}N_{j},$
and by (4.7) and (4.9), it is immediate to see that
$N_{0}\leq c_{2}\,c_{3}\,N^{(d-1)/(4d)}.$ (4.10)
Last, consider the constant depending on $u$ defined as
${\rm M}(u)=\max_{|{n}|\geq 1}|{n}|^{-d}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi,$
and in particular, notice that this quantity is indeed finite. Finally, for $11\leq j\leq K+$ 1, we assign a uniform lattice $\mathcal{P}_{N_{j}}$ to every $N_{j}$ as in (4.6), and by proceeding as in the first part of the proof, we get the estimate
$\mathcal{D}_{2}(u;\mathcal{P}_{N_{j}})\leq\min\left(c_{1}\mathcal{J}(u)N_{j}^{(d-1)/d},\,c_{4}M(u)n_{0}^{d-1}\right),$ (4.11)
where the minimum involved depends on whether or not it holds $N_{j}^{1/d}\geq n_{0}$. On the other hand, for every set of $N_{0}$ points $\mathcal{P}_{N_{0}}\subset\mathbb{T}^{d}$, by the inequality in (4.10), we get
$\begin{split}\mathcal{D}_{2}(u;\mathcal{P}_{N_{0}})&=\sum_{{n}\in\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{{p}\in\mathcal{P}_{N_{0}}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\\ &\leq c_{5}{\rm M}(u)N^{(d-1)/(2d)}.\end{split}$ (4.12)
Therefore, by (4.11) and (4.12), we get
$\begin{split}\mathcal{D}_{2}(u;\mathcal{P}_{N})&=\sum_{{n}\in\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{j=0}^{K+1}\sum_{{p}\in\mathcal{P}_{N_{j}}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\\ &\leq(K+2)\sum_{j=0}^{K+1}\sum_{{n}\in\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{{p}\in\mathcal{P}_{N_{j}}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\\ &\leq(K+2){\rm M}(u)\left(c_{4}(K+1)n_{0}^{d-1}+c_{5}N^{(d-1)/(2d)}\right)+c_{1}(K+2)\mathcal{J}(u)\sum_{j=1}^{K+1}N_{j}^{(d-1)/d}\\ &\leq c_{5}\,{\rm M}(u)\,N^{(d-1)/(2d)}+C_{d}\,\mathcal{J}(u)\,N^{(d-1)/d},\end{split}$
[…]
We need a technical lemma to obtain the lower bound. We briefly present a celebrated result of Cassels and Montgomery; the proof is inspired by Siegel’s analytic proof of Minkowski’s convex body theorem.
Lemma 4.4 (Cassels-Montgomery).
For every origin-symmetric convex body $\Omega\subset\mathbb{R}^{d}$, and for every finite set of points $\{{p}_{j}\}_{j=1}^{N}\subset\mathbb{T}^{d}$, it holds
$\sum_{{m}\in\Omega\cap\mathbb{Z}_{*}^{d}}\bigg{|}\sum_{j=1}^{N}e^{2\pi i{m}\cdot{p}_{j}}\bigg{|}^{2}\geq 2^{-d}|\Omega|N-N^{2}.$
Proof.
Consider the auxiliary sets $A_{\Omega}({x})\subset\mathbb{Z}^{d}$ defined by
$A_{\Omega}({x})=\left({x}+\Omega/2\right)\cap\mathbb{Z}^{d}.$
Notice that
$\int_{\mathbb{T}^{d}}\#A_{\Omega}({x})\,\,\mathrm{d}{x}=\int_{\mathbb{T}^{d}}\sum_{{n}\in\mathbb{Z}^{d}}\mathds{1}_{\Omega/2}({n}-{x})\,\,\mathrm{d}{x}=\int_{\mathbb{R}^{d}}\mathds{1}_{\Omega/2}({x})\,\,\mathrm{d}{x}=2^{-d}|\Omega|,$
and therefore we find an $\tilde{{x}}\in[0,1)^{d}$ such that #$\#A_{\Omega}(\tilde{{x}})\geq 2^{-d}|\Omega|$. Now, consider the non-negative trigonometric polynomial
$T({y})=\frac{1}{\#A_{\Omega}(\tilde{{x}})}\bigg{|}\sum_{{n}\in A_{\Omega}(\tilde{{x}})}e^{2\pi i{n}\cdot{y}}\bigg{|}^{2}=\frac{1}{\#A_{\Omega}(\tilde{{x}})}\sum_{{n},{m}\in A_{\Omega}(\tilde{{x}})}e^{2\pi i({n}-{m})\cdot{y}},$
and notice that the function of its Fourier coefficients $\widehat{T}\colon\mathbb{Z}^{d}\to\mathbb{R}$ is non-negative as well and, since ${n},{m}\in A_{\Omega}(\tilde{{x}})$ imply $({n}-{m})\in\Omega$, then its support is contained in $\Omega$. Further, observe that we have
$T(0)=\#A_{\Omega}(\tilde{{x}})\geq 2^{-d}|\Omega|,$
moreover, it is not difficult to notice that for all ${n}\in\mathbb{Z}^{d}$, it holds
[…]
Therefore, it follows that
$\begin{split}\sum_{{n}\in\Omega\cap\mathbb{Z}^{d}}\left|\sum_{j=1}^{N}e^{2\pi i{n}\cdot{p}_{j}}\right|^{2}&\geq\sum_{{n}\in\Omega\cap\mathbb{Z}^{d}}\widehat{T}({n})\left|\sum_{j=1}^{N}e^{2\pi i{n}\cdot{p}_{j}}\right|^{2}\\ &=\sum_{j=1}^{N}\sum_{\ell=1}^{N}\sum_{{n}\in\Omega\cap\mathbb{Z}^{d}}\widehat{T}({n})\,e^{2\pi i{n}\cdot({p}_{j}-{p}_{\ell})}\\ &=\sum_{j=1}^{N}\sum_{\ell=1}^{N}T({p}_{j}-{p}_{\ell})\geq N2^{-d}|\Omega|,\end{split}$
and the conclusion easily follows ∎
Last, we show Theorem 1.6 through the previous lemma.
Proof of Theorem 1.6.
We will write $c_{i}=c_{i}(d)$ with $i\in\mathbb{N}$ to intend generic dimensional constants throughout the proof. By (4.4), there exists $n_{0}\in\mathbb{N}$ be such that for every ${n}\geq n_{0}$, it holds
$\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\geq\frac{{|{n}|^{d}\,\mathcal{J}}(u)}{4d\pi^{2}}.$ (4.13)
Hence, consider
${\rm m}(u)=\min_{1\leq|{n}|\leq n_{0}}|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi.$
First, we show the claim holds for all $u$ such that ${\rm m}(u)>0$, and in particular notice that ${\rm m}(u)=0$ if and only if $\hat{u}$ is zero in $B_{1}(\mathbf{0})$. We prove the lower bound for all positive integers $N$ such that
$\frac{\mathcal{J}(u)}{4d\pi^{2}}N^{-(d+1)/d}\leq{\rm m}(u),$
and for every set of $N$ points $\mathcal{P}_{N}\subset\mathbb{T}^{d}$. Indeed, within this assumption, for all $n\in B_{N^{1/d}}(\mathbf{0})$ it holds
$|{n}|^{-2d-1}\int_{B_{|{n}|}(\mathbf{0})}|\xi|^{d+1}|\hat{u}|^{2}(\xi)\,\,\mathrm{d}\xi\geq\frac{\mathcal{J}(u)}{4d\pi^{2}}N^{-(d+1)/d}.$
Therefore, by applying the latter inequality in (4.3), we get that for a dimensional constant $\kappa\geq 1$ to be chosen later, it holds
$\mathcal{D}_{2}(u;\mathcal{P}_{N})\geq\sum_{{n}\in\mathbb{Z}_{*}^{d}\cap B_{(\kappa N)^{1/d}}(\mathbf{0})}\bigg{|}\sum_{{p}\in\mathcal{P}_{N}}e^{2\pi i{p}\cdot{n}}\bigg{|}^{2}\frac{\mathcal{J}(u)}{4d\pi^{2}}(\kappa N)^{-(d+1)/d}.$
Finally, we can apply the previous lemma and get
$\mathcal{D}_{2}(u;\mathcal{P}_{N})\geq c_{1}\,\mathcal{J}(u)\,(\kappa N)^{-(d+1)/d}\left(|B_{1}(\mathbf{0})|\kappa N^{2}-N^{2}\right),$
and by setting $\kappa=2|B_{1}(\mathbf{0})|^{-1}$, it follows that
$\mathcal{D}_{2}(u;\mathcal{P}_{N})\geq c_{2}\,\mathcal{J}(u)\,N^{(d-1)/d}.$
Now, by contradiction, suppose there exists $u\in{\rm BV}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$ be such that $\hat{u}$ is zero in $B_{1}(\mathbf{0})$, and such that there exists an increasing sequence of integers $\{N_{j}\}_{j\in\mathbb{N}}$, with corresponding sets of $N_{j}$ points $\mathcal{P}_{N_{j}}\subset\mathbb{T}^{d}$, such that
$\mathcal{D}_{2}(u;\mathcal{P}_{N_{j}})\leq\frac{c_{2}}{4}\,\mathcal{J}(u)\,N_{j}^{(d-1)/d}.$ (4.14)
Now, notice that for a continuous function $v\in{\rm BV}(\mathbb{R}^{d})\cap L^{\infty}(\mathbb{R}^{d})$ such that $\hat{v}(\mathbf{0})>0$, the claim holds with $\mathcal{J}(v)=0$. Similarly, the claim would also hold for $u+v$ since
$\mathcal{F}\{u+v\}(\mathbf{0})=\hat{v}(\mathbf{0})>0,$
and therefore, since $\mathcal{J}(u+v)=\mathcal{J}(u)$, we would get
$\liminf_{N\to+\infty}N^{(1-d)/d}\inf_{\mathcal{P}_{N}\subset\mathbb{T}^{d}}\mathcal{D}_{2}(u+v;\mathcal{P}_{N})\geq c_{2}\,\mathcal{J}(u).$
Last, we notice that by the triangular inequality and (4.14), for every $j\in\mathbb{N}$ it holds
$\begin{split}\mathcal{D}_{2}(u+v;\mathcal{P}_{N_{j}})&\leq 2\mathcal{D}_{2}(u;\mathcal{P}_{N_{j}})+2\mathcal{D}_{2}(v;\mathcal{P}_{N_{j}})\\ &\leq\frac{c_{2}}{2}\,\mathcal{J}(u)\,N_{j}^{(d-1)/d}+o\big{(}N_{j}^{(d-1)/d}\big{)},\end{split}$
and this is absurd. Therefore, the claim holds with $C_{d}=c_{2}/4$.